Integration of x square divided by three minus cosx

Grade:12th pass

1 Answers

Arun
4 years ago
Dear student
 
\int (x^2 /3 ) - cos x = (x^3/9) - sin x + constant
 
Hope it helps
In case of any difficulty, please feel free to ask again.
 
Thanks & Regards
Arun